OpenStudy (snowcrystal):

Can someone please help me? :( I have no clue how to do this. I will fan and medal.

OpenStudy (snowcrystal):

(2*x*2x+5) + (2*2x+5*x+3) + (2*x*x+3)

OpenStudy (aayushi.somani):

what does * means? is it multilplication?

OpenStudy (snowcrystal):

yah the * is a multiplacation sign cuz i have x as a varriable

OpenStudy (aayushi.somani):

4x^2+5+4x+5x+3+2x^2+3 6x^2+9x+6 now solve the given equation..! can u do so.?

OpenStudy (snowcrystal):

do you add both of those problems together?

OpenStudy (aayushi.somani):

there is only 1 problem i can see,,!! @SnowCrystal

OpenStudy (***[isuru]***):

\[( 2 \times x \times 2x + 5) = 4x^{2} + 5\] \[( 2 \times 2x + 5 \times x + 3) = 4x + 5x + 3 = 9x + 3\] \[( 2 \times x \times x + 3 ) = 2x^{2} + 3\] now add them together :)

OpenStudy (***[isuru]***):

\[4x^{2} + 5 + 9x + 3 + 2x^{2} + 3 = ?\]

OpenStudy (aayushi.somani):

yea m mistaken 6x^2+9x+8

OpenStudy (snowcrystal):

umm that would be.....11+9x+6x^2 right?

OpenStudy (***[isuru]***):

yep !

OpenStudy (snowcrystal):

would th

OpenStudy (snowcrystal):

my finial answer?

OpenStudy (aayushi.somani):

oh yea..!! i missed the 3..!! :P

OpenStudy (snowcrystal):

i hate computers it is typing on its own

OpenStudy (***[isuru]***):

yes its ur final answer :)

OpenStudy (snowcrystal):

thanks guys :)

OpenStudy (***[isuru]***):

u r welcome!

OpenStudy (aayushi.somani):

welcum..!